OpenStudy (dmezzullo):

Choose the equation that could be used to find two consecutive integers whose sum is 69. n + (n + 2) = 69 n + (n + 1) = 69 n − (n + 1) = 69 n + 2n = 69

OpenStudy (dmezzullo):

@jazy

OpenStudy (nubeer):

first integer =n second consecutive integer = n+1 sum of both numbers = 69 n+(n+1) = 69

OpenStudy (anonymous):

let n=23 2n= 2*23=46 n+2n=69

OpenStudy (nubeer):

it says consecutive numbers. not any numbers